OpenStudy (anonymous):

what is the LCM of 14x^2+28 and 7x^2+42x+56 what is the LCM of 14x^2+28 and 7x^2+42x+56 @Mathematics

OpenStudy (anonymous):

i think i have the 14x^2+28 factored right is it 14x(x+2)=2x7x(x)x(x+2)

OpenStudy (anonymous):

first you want to find their GCD. Then use the formula: \[LCM(a,b)=\frac{a\cdot b}{GCD(a.b)}\]

jhonyy9 (jhonyy9):

14(x2 +2)= 7(x2 +6x +8)=7(x+4)(x+2)

OpenStudy (anonymous):

\[14x^2+28=14(x^2+2)\]

OpenStudy (anonymous):

Looks like their GCD is only 7. So the LCM would be:\[\frac{(14x^2+28)(7x^2+42x+56)}{7}\]

OpenStudy (anonymous):

the GCD is 7, not the LCM. The LCM will be bigger than the two polynomials.

OpenStudy (anonymous):

so after i factor the second is where i get lost

jhonyy9 (jhonyy9):

14(x2 +2)=(2(7(x+2) on exponent2) - 4x) and hence LCM will be 7(x+2)

OpenStudy (anonymous):

you factor the first and get:\[14(x^2+2)=2\cdot7(x^2+2)\] factoring the second gives:\[7x^2+42x+56=7(x^2+6x+8)=7(x+2)(x+4)\] Their only common factor is 7, that is the GCD. Then use the formula i posted above.

OpenStudy (anonymous):

somehow i came up with 14x(x+2)(x+4)

OpenStudy (anonymous):

there is no way the LCM can be 7(x+2). what would you multiply 14x^2+28 by to get 7x+14?

OpenStudy (anonymous):

yes joe u were ryt LCM is 14*(x^2+2)*(x+2)*(x+4)

OpenStudy (anonymous):

but how i got that answer i have no idea lol

OpenStudy (anonymous):

sheg has the correct answer.

OpenStudy (anonymous):

thanks all!

OpenStudy (anonymous):

Think of how you would do this with numbers. What is the LCM of 6 and 9? its 18 because 6*3 is 18, and 9*2 is 18, and 18 is the smallest common multiple between the two. Using the formula above, the GCD of 6 and 9 is 3, so the LCM = 6*9/3=18.